# adversarial attack DB Diffusion Buffer

### v2

@@ 3.3 Projected Gradient Descent with an $\ell_{2}$ budget
10 10 $1.84\pm 0.43$ $3\pm 0.04 0.11±0$$\pm 0.11 1.1$$15\pm 0.36$ $3.17\pm 0.64$ $0.77\pm 0.06$ $10.26\pm 5.32$
Table 1: Targeted-attack results on 100 EARS-WHAM-v2 pairs. Left block of metrics reports attack success for the enhanced signal $\hat{S}^{Y^{\mathrm{user}}+\delta}$ against the target $S^{\mathrm{attacker}}$; right block reports perturbation impact for $(Y^{\mathrm{user}}{+}\delta)$ against $Y^{\mathrm{user}}$ and the SNR of $\delta$ w.r.t. $Y^{\mathrm{user}}$ (higher SNR = quieter perturbation).
@@ 4.3 Evaluation metrics
Table 2: Ablation of diffusion SE on the same 100 pairs. Unless specified, the sampler uses N=25N{=}25 reverse steps and stochastic sampling; Fixed-noise denotes a frozen noise path (Sec.  2.2). The baseline model is trained with $\sigma_{\mathrm{max}}{=}0.5$; rows that list $\sigma_{\mathrm{max}}$ are separately trained models.
@@ 7 Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 545210893, 498394658. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project f102ac. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. The authors would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.

### v3

@@ 3.3 Projected Gradient Descent with an $\ell_{2}$ budget
10 10 $1.84\pm 0.43$ $3\pm 0.04 0.11±0$$\pm 0.11 1.1$$15\pm 0.36$ $\bm{3.17\pm 0.64}$ $\bm{0.77\pm 0.06}$ $\bm{10.26\pm 5.32}$
Table 1: Targeted-attack results on 100 EARS-WHAM-v2 pairs. Left block of metrics reports attack success for the enhanced signal $\hat{S}^{Y^{\mathrm{user}}+\delta}$ against the target $S^{\mathrm{attacker}}$; right block reports perturbation impact for $(Y^{\mathrm{user}}{+}\delta)$ against $Y^{\mathrm{user}}$ and the SNR of $\delta$ w.r.t. $Y^{\mathrm{user}}$ (higher SNR = quieter perturbation). Cell backgrounds show a per-metric min-max colormap (respecting $\uparrow/\downarrow$), and the best value within each block is bolded. Values are reported as mean $\pm$ standard deviation.
@@ 4.3 Evaluation metrics
Table 2: Ablation of an unconstrained adversarial attack on diffusion SE ($\lambda{=}-,\varepsilon{=}\infty$). Unless specified, the sampler uses N=25N{=}25 reverse steps and stochastic sampling; Fixed-noise denotes a frozen noise path (Sec. 2.2). The baseline model is trained with $\sigma_{\mathrm{max}}{=}0.5$; rows that list $\sigma_{\mathrm{max}}$ are separately trained models. Cell backgrounds show a min–max colormap per metric using the same color ranges as Table 1
@@ 7 Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 545210893, 498394658. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project f102ac. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. This research was partially funded by VolkswagenStiftung Niedersächsisches Vorab - ZN4704 and the Daimler and Benz Foundation under the grant Ladenburger Kolleg, Project KonCheck. We acknowledge funding by the German Federal Ministry of Research, Technology and Space (BMFTR) under grant agreement No. 16IS24072B (COMFORT), SisWiss (16KIS2330), and AIgenCY (16KIS2012). The authors would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.